there is 116.9 g Mg 225.2 g Cr and 61.60 g O write the empirical formula. I end of with 1.25 for Mg 1.125 for Cr and 1 for O . Since these are not whole numbers I am stumped
1
Expert's answer
2020-12-21T03:39:23-0500

moles of Mg = 116.9/24 = 4.87moles

moles of Cr = 225.2/52 = 4.33 moles

moles of oxygen = 61.60/16 = 3.85 moles


ratio is 4.87: 4.33 : 3.85

in more shorter form 1.26 : 1.124 : 1

we can write as 12.6 : 11.24 : 10

12.6 nearly equal to 13

11.24 nearly equal to 11

so empirical formula is Mg13Cr11O10

It is not possible to have neutral molecule with this empirical formula